[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Sep 25, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Sep 25 16:39:59 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt S.

----------

TAFB:
1. GFE Issue: there were a couple of instances of Jacobson SWELL formatter  
failure over the weekend. According to GFE focal point Jeff Lewitsky, no  
need to log failure of this formatter in the shift log, just use the  
official OFF formatter instead.

2. Problem reported on Saturday, 9/23 with GOES-16 imagery not loading on  
D2D on LX3. Received "Error loading style rules" message. Rebooting  
workstation did not solve the problem. Experienced same problem with the  
other TAFB workstations but problem did not extend to the HSU workstations.  
Chris Mello opened NCF trouble ticket # 52711 for this outage. NCF followed  
up on Sunday, 9/24, problem had resolved.
HSU:
1. Cable TV is not working in HSU, even though it is working in TAFB and  
the media room.

2. Website failed when special advisory was issued. Problem was related to  
using 0630 UTC as the issuing time. Dave Zelinsky manually fixed problem on  
www01 server and restarted halo.

3. Brief hiccup between WOC and Amazon Web Services caused a few storm  
table thumbnails and Maria archive to be missing/Forbidden on public site.  
Problem not observed on local server. Auto-resolved after a few minutes.

4. EPAC ATCF workstation and N-AWIPS box SAWFISH hung up around 09:00 UTC.  
To prevent/minimize these kind of issues, TSB recommends rebooting or at  
least logging out of all ATCF and N-AWIPS workstations at least once this  
week.
Other items:
1. NHC submitted their GOES-16 meso-2 extension request for
Hurricane Maria, to now end at 9/25/ 2100Z.
2. CWD for Hurricane Maria to run from 9/24 21:45Z - 9/28 00:00Z
3. AMSU size analyses (wind radii) were only showing up for one storm per  
basin in N-AWIPS. Matt S. fixed this.
4. There was a problem viewing inundation graphic on the nhc-vm-webdev01  
server, restarting Halo resolved (Dave Zelinsky).

NHC RFCs:
RFC 252 - Upgrade ATCF FTP scripts to use rsync instead of scp; implement  
9/27 at 14:00Z, postponed due to multiple CWDs.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170925/cc518213/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list